A firm has a market capitalization (market value of equity) of $11 Billion and net debt of $3 Billion. Calculate the weight of equity in the firm's weighted average cost of capital (WACC) calculation. 
Answer #1

total value of firm = value of equity + value of debt

= 11 + 3

= 14 billion

weight of equity = value of equity/total value of firm

= 11/14

= 78.57%
Answer #2

ANSWER :




C = E + D = 11 + 3 = 14 billion dollars.


Weight of equity in the capital = E/C = 11 / 14 = 0.7857 = 78.57% (ANSWER).

answered by: Tulsiram Garg